A corner lot has dimensions 25 by 40 yards. The city plans to take a strip of uniform width along the two sides bordering the streets to widen these roads. How wide should the strip be if the remainder of the lot is to have an area of 844 square yards? (Round your answer to two decimal places.)
Answer by jorel555(1290)

Let w be the width of the strip. Then your new length would be 40-2w, and your new width would be 25-2w. So:
(40-2w)(25-2w)=844
1000-130w+4w²=844
4w²-130w+156=0
Using the quadratic formula, we get w=31.2520831887 or 1.24791681132.
Throwing out the larger result, we get the width of the strip to be approx. 1.25 yds. ☺☺☺☺